OpenStudy (anonymous):

The perimeter of a rectangle is 36 ft. The length is 6 ft longer than the width. Find the dimensions. write a system of linear equations and solve the resulting system. let x be the length and y be the width.

OpenStudy (tkhunny):

Please express the perimeter of a rectangle in terms of its length (x) and width (y).

OpenStudy (anonymous):

2x+2y=36

OpenStudy (tkhunny):

Very good. Now, translate this: " The length is 6 ft longer than the width. " In terms of x and y.

OpenStudy (anonymous):

y=x+6

OpenStudy (tkhunny):

I really hate these variable names. How do you remember which is length and which is width? It would be so much easier if L = Length and W = Width. Anyway, we have x = length. This is the side of greater measure. We also have y = width. If x is already greater than y, by definition, how does x+6 compare to y?

OpenStudy (anonymous):

lower

OpenStudy (tkhunny):

x > y -- x is greater x + 6 >> y -- x+6 is WAY greater! So, y = x+6 is no good. You should have x = y+6. y is shorter. You must increase the shorter by 6 to make them equal.

OpenStudy (anonymous):

w=6 and L=12

OpenStudy (tkhunny):

That is indeed one 6 greater than the other. Does it work in the perimeter formula?

OpenStudy (anonymous):

thats the answer i got for it all i get confused with this

OpenStudy (tkhunny):

We have: 2x + 2y = 36 and x = y+6 Can you find a simultaneous solution?